OpenStudy (anonymous):

y = 1/2x - 6 2x + 6y = 19 Can you show me both Elimination and Substitution method?

OpenStudy (johnweldon1993):

We'll start with the substitution method...you know what y = since y = 1/2x - 6 so just substitute that into the second equation wherever you see a 'y' 2x + 6(1/2x - 6) = 19 can you solve this for 'x' ?

OpenStudy (anonymous):

2x + (3x - 36) = 19 Am I doing this correctly..?

OpenStudy (johnweldon1993):

You are indeed :)

OpenStudy (anonymous):

So then do I add 36 onto the -36 and 19?

OpenStudy (johnweldon1993):

Yes add 36 to both sides... so then you will have 2x + 3x = 55 And from there...?

OpenStudy (anonymous):

6x = 55

OpenStudy (anonymous):

wait 5x I mean

OpenStudy (anonymous):

So x would equal 11

OpenStudy (johnweldon1993):

that is correct! now that you know what x equals...go back into either original equation and plug in x = 11...and solve for y... lets do the second one... 2x + 6y = 19 becomes 22 + 6y = 19 Solving for y we get...?

OpenStudy (anonymous):

Well you would subtract 22 on both sides, so it would be 6y = -3 ...?

OpenStudy (johnweldon1993):

Right....so what does y = ? what is -3 / 6 ?

OpenStudy (anonymous):

-0.5 ?

OpenStudy (johnweldon1993):

Right! lets keep it as a fraction though...so x = 11 y = -1/2 Okay ready for the elimination method?

OpenStudy (anonymous):

Yeah, I am. I didn't realize how easy that was.

OpenStudy (johnweldon1993):

lol yeah they are simple :) ...well actually...do you know the elimination method? if you do....do this out for me since you know the answers you SHOULD get...want to do this one and let me know what you get? or should we do it out here?

OpenStudy (anonymous):

Actually the elimination method is more difficult for me. Can we work it out..?

OpenStudy (johnweldon1993):

Absolutely :)

OpenStudy (anonymous):

Actually, I have to go eat. Can we continue this in like 20 minutes..?

OpenStudy (johnweldon1993):

haha sure I'll still be around :)

OpenStudy (anonymous):

Alright thanks ^-^

OpenStudy (anonymous):

Back c:

OpenStudy (johnweldon1993):

How was dinner? lol

OpenStudy (anonymous):

Oh delishh c: Thank you ^-^

OpenStudy (johnweldon1993):

lol of course :) so okay ready for the elimination method?

OpenStudy (anonymous):

Yes, sir c:

OpenStudy (johnweldon1993):

Hang on it's a lot to type lol

OpenStudy (anonymous):

Haha no worries, take your time c:

OpenStudy (johnweldon1993):

y = 1/2x - 6 2x + 6y = 19 Lets line up the variables... y = 1/2x - 6 6y = -2x + 19 Okay that looks better... so now we want to eliminate one of these variables...doesn't matter which one...lets choose 'x'...so multiply the top equation by 4 4y = 2x - 24 6y =-2x +19 Now we want to combine (add) these 2 equations... 4y = 2x - 24 + 6y = -2x + 19 ---------------- 4y + 6y = 10y right? 2x - 2x = 0 right? and -24 + 19 = -5 right? so we have 10y = -5 y = what ?

OpenStudy (johnweldon1993):

Tell me if there is anything you don't get...

OpenStudy (anonymous):

So would y equal -1/2? Or did I divide it wrong?

OpenStudy (johnweldon1993):

No you got it right :) make sense? what did we find y to equal in the substitution method? -1/2 ...so we are on par here :)

OpenStudy (anonymous):

Yeah it makes sense, oh I see.

OpenStudy (johnweldon1993):

Alright...so to solve for 'x'...what would we do? any idea?

OpenStudy (anonymous):

Don't we replace the y, with the answer we got?

OpenStudy (johnweldon1993):

You're thinking of the substitution method again :P nope we need to do the entire process over again and solve for 'x' this time lol

OpenStudy (anonymous):

Oh lord /.\ Bring it on xD

OpenStudy (johnweldon1993):

lol Okay... y = 1/2x - 6 6y = -2x + 19 Now you tell me...how would we eliminate the 'y' variable this time?

OpenStudy (anonymous):

Multiply by 6 on the bottom equation, correct?

OpenStudy (johnweldon1993):

You mean the top equation right? if you did the bottom there would be y = 1/2x - 6 36y = -12x + 114 not quite :)

OpenStudy (johnweldon1993):

Wow this is not adding up for me...okay hang on lol...

OpenStudy (anonymous):

Haha, yeah I was looking at it

OpenStudy (johnweldon1993):

alright so y - 1/2x = -6 6y + 2x = 19 multiply the top equation by -6 -6y + 3x = 36 6y + 2x = 19 add the 2 equations... -6y + 3x = 36 + 6y + 2x = 19 --------------- 5x = 55 ahh there we go...don't know why I kept messing up lol

OpenStudy (anonymous):

So yeah, then x would equal 11 c:

OpenStudy (johnweldon1993):

yeah lol....apparently on paper I can't work correctly lol... but yes x would = 11 lol and there we have both methods...obviously if you can....substitute haha :)

OpenStudy (anonymous):

Ah god,

OpenStudy (anonymous):

You helped me so much

OpenStudy (johnweldon1993):

Glad to do so :) hope everything was understandable :)

OpenStudy (anonymous):

Very c: Thank you so much!

OpenStudy (johnweldon1993):

Anytime :)

OpenStudy (anonymous):

I have another one, but I might finish it tomorrow xD And it seems more understandable so c:

OpenStudy (anonymous):

But I have to go, thanks again!

OpenStudy (johnweldon1993):

Alright if you can't get it feel free to message me on here :) have a good night :)